A block of ice at 0oC whose mass initially is m = 42.1 kg slides along a horizontal surface, starting at a speed v0 = 2.35 m/s and finally coming to rest after traveling a distance d = 2.93 m. Compute the mass of ice melted as a result of the friction between the block and the surface. (Assume that all the heat generated owing to friction goes into the block of ice.)
Answer #1

Initial KE = 0.5*0.0421*2.35^2 = 0.1162 J
Final KE = 0
Heat = KE = 0.1162 J = mL = m*334000
m = 0.348 mg